OpenStudy (anonymous):

How much pure water must be mixed with 10 liters of a 25% acid solution to reduce it to a 10% acid solution?

hero (hero):

Liters Equation: x + 10 = y Water Equation: 1x + .75(10) = .90y x = x y - 10 = .90y - 7.5 y - .90y = 10 - 7.5 .10y = 2.5 y= 2.5/.10 y = 25 x = 15 Therefore `15 liters` of pure water was added, which produced a total of 25 liters of a 10% solution.